OpenStudy (jacky90):

HELP Petrolyn motor oil is a combination of natural oil and synthetic oil. It contains 5 liters of natural oil for every 7 liters of synthetic oil. In order to make 624 liters of Petrolyn oil, how many liters of natural oil are needed?

OpenStudy (kikimarie6779):

the answer is 260 to 364

OpenStudy (aaronq):

You need to know the composition of natural oil in petrolyn oil. \(\sf fractional~composition~of~Natural~oil=\dfrac{L~of~Natural~oil}{L~of~Natural~oil+L~of~Synthetic~oil}=\dfrac{5}{5+7}\) Now, you have 624 L of Petrolyn oil but want to know the L of Natural oil, so you multiply the fractional composition of Natural oil by the total amount of oil. \(\sf L~of~Natural~oil=\dfrac{5}{12}*624~L\)

OpenStudy (jacky90):

Thanks =)

OpenStudy (aaronq):

no probs !